OpenStudy (anonymous):

In a mixture of concrete, there are two lb of cement mix for each pound of gravel. If the mixture contains a total of 120 lb of these two ingredients, how many pounds of gravel are there?

OpenStudy (anonymous):

The total pounds in a simple mixture is 3, where 2 are cement and 1 is gravel. So for cement: \[ \frac{2}{3}=\frac c{120} \]And for gravel: \[ \frac 13=\frac g{120} \]

OpenStudy (anonymous):

What do I do next? To solve

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

solve for each

OpenStudy (anonymous):

I haven't crossed multiplied since freshman year. So, I got 80 for the cement and 40 for gravel. What do I do next?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%2F3%3Dc%2F120%2C+1%2F3%3Dg%2F120 That is correct.

OpenStudy (anonymous):

That's it?

OpenStudy (anonymous):

Yes. It just asks for the pounds of gravel.